OpenStudy (anonymous):

The label on a juice carton says there are 50 calories in a 4-ounce serving of juice. If Arthur drinks 12 ounces of juice, how many calories will he consume? A. 150 calories B. 58 calories C. 48 calories D. 200 calories

OpenStudy (girllover123):

@Donblue22

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@taymiyyah Please do not give direct answers. They are against OpenStudy's Code of Conduct. If you keep giving them, you might get suspended. Please refrain from giving direct answers in the future. Thank you. :) If you do not know the rules, then you can read them here: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

mak a proprtion aagain

OpenStudy (anonymous):

50/4 = 12/x

OpenStudy (anonymous):

Set up a proportion... ._.

OpenStudy (anonymous):

corss multiply 50 * x = ? 4 * 12 = ?

OpenStudy (anonymous):

I thought that's 50/4 = x/12...?

OpenStudy (anonymous):

oh, that's right..i read it wrong

OpenStudy (anonymous):

Hehe

OpenStudy (anonymous):

\(\dfrac{50}{4}\) = \(\dfrac{x}{12}\) 50 * 12 4 * x

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Multiply that up there

OpenStudy (anonymous):

k so 50 *12 and 4*x so 50 *12=600 and 4*x= so..... ?

OpenStudy (anonymous):

4 * x = 4x

OpenStudy (anonymous):

So you have 4x = 600. Divide 600 by 4.

OpenStudy (anonymous):

so 150

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

k thx